Circular No. 6646

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE The High-Z Supernova Search Team (cf. IAUC 6160, plus J. Tonry,

University of Hawaii; A. V. Filippenko, University of California, Berkeley; R. Gilliland, Space Telescope Science Institute; and D. Reiss and A. Diercks, University of Washington) reports the discovery of eleven supernovae on CCD images taken at the Canada-France-Hawaii Telescope (CFHT) by P. Garnavich and J. Jensen. The supernovae were identified by B. Schmidt and J. Tonry by comparing images taken on Apr. 29 and 30 UT with those obtained on Apr. 4 and 5.

SN  1997 UT  α_{2000}  δ_{2000}  I  z type
1997ca Apr. 29 14^{h}16^{m}40.^{s}24 +52^{o}31′59′′.2 22.6 0.52 Ia
1997cb Apr. 29 14 17 55.41 +52 19 00.6 23.7 0.28 II?
1997cc Apr. 29 14 16 33.59 +52 23 37.7 23.5
1997cd Apr. 29 17 06 17.07 +44 04 34.0 22.9 0.51 Ia
1997ce Apr. 29 17 07 48.26 +44 01 26.2 23.3 0.44 Ia
1997cf Apr. 29 17 06 43.08 +43 54 14.0 23.7
1997cg Apr. 29 17 06 46.96 +43 49 14.1 23.5 ?  ?
1997ch Apr. 29 12 37 28.92 +62 06 42.8 23.5
1997ci Apr. 29 12 37 13.41 +62 05 20.5 23.5
1997cj Apr. 29 12 37 04.27 +62 26 24.9 22.8 0.50 Ia
1997ck Apr. 29 16 53 00.27 +35 03 46.4 24 0.97 ?

The supernovae are within 3′′ of their host galaxies, except SN 1997ca, which has no detectable host. For the supernovae with indicated redshift and type, spectra were obtained with the Multiple-Mirror Telescope on May 2–3 by P. Challis (reductions by B. Leibundgut) and with the Keck telescope on May 3–4 by B. Schmidt, E. Hu, and L. Cowie. SNe 1997cc and 1997ch have very blue V − I colors and may be type-II supernovae. With z ∼ 1, SN 1997ck is the most distant supernova yet observed. Finder charts may be requested by sending electronic mail to pgarnavich@cfa.harvard.edu.

SUPERNOVA 1997bq IN NGC 3147 Magnitude estimates: May 2.8 UT, 15.1 (S. Pesci, Milan, Italy, visual);

2.93, 15.0 (K. Sárneczky and L. Kiss, Szeged, Hungary, unfiltered CCD).

1997 May 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green